# The Combinatorics of Motzkin Polyominoes

### v1

@@ Abstract.
A word $w=w_{1}\cdots w_{n}$ over the set of positive integers is a Motzkin word whenever $w_{1}=\texttt{1}$, $11\leq w_{k}\leq w_{k-1}+$ 1, and $w_{k-1}\neq w_{k}$ for $k=2,\dots,n$. It can be associated to a $n$-column Motzkin polyomino whose $i$-th column contains $w_{i}$ cells, and all columns are bottom-justified. We reveal bijective connections between Motzkin paths, restricted Catalan words, primitive Łukasiewicz paths, and Motzkin polyominoes. Using the aforementioned bijections together with classical one-to-one correspondence with Dyck paths avoiding $UDU$s, we provide generating functions with respect to the length, area, semiperimeter, value of the last symbol, and number of interior points of Motzkin polyominoes. We give asymptotics and close expressions for the total area, total semiperimeter, sum of the last symbol values, and total number of interior points over all Motzkin polyominoes of a given length. We also present and prove an engaging trinomial relation concerning the number of cells lying at different levels and first terms of the expanded $(1+x+x^{2})^{n}$.
@@ 1. Introduction
These sequences have a similar behavior since, almost always, they appear together, and a large number of similar types of combinatorial classes are enumerated by these numbers. For Catalan numbers, we refer to the catalog of Catalan combinatorial classes compiled by Stanley [26]. Similarly, there are several combinatorial classes where Motzkin family arises. For example, the rooted plane trees are a classical Catalan family, but if we add loops, yields a Motzkin family [13]. Additional examples of Motzkin families can be found in [3, 7] and the references therein.
@@
In the following of this work, we focus on the class of Motzkin polyominoes, and we investigate enumerative problems about several statistics on these objects.
@@
Motivation and outline of the paper. The motivation of this work is to provide enumerative results on Motzkin words according to several parameters (area, semiperimeter, and number of interior points) defined on the associated polyominoes. Except for the first section where we exhibit bijections between Motzkin words and other classical combinatorial classes counted by the Motzkin numbers, all other results are obtained algebraically by considering functional equations for multivariate generating functions. In Section 2 we give three bijections between Motzkin words and three known combinatorial classes counted by the Motzkin numbers. In Section 3, we focus on the two statistics of the area and the semiperimeter. We provide the trivariate generating function, where the coefficient of $x^{n}p^{k}q^{\ell}$ is the number of Motzkin polyominoes $P(w)$ of length $n$, and satisfying ${\texttt{sper}}(w)=k$ and ${\texttt{area}}(w)=\ell$. From this, we deduce the generating function for the total sum of the last symbol over all Motzkin words of length $n$, and we give asymptotic approximation for the coefficient of $x^{n}$. In Section 4, we focus on the semiperimeter statistic. We give the generating function for nonempty Motzkin polyominoes with respect to the length and the semiperimeter. We exhibit a bijection between Motzkin polyominoes of length $n$ with semiperimeter $2n-k$, and Motzkin paths of length $n-1$ with exactly $k$ up steps. We deduce a close expression for the total semiperimeter over all Motzkin polyominoes of length $n$. In Section 5, we make a similar study as in Section 4 by considering the area instead of the semiperimeter. Finally, Section 6 is dedicated to the statistic of the number of interior points of the Motzkin polyominoes, and we exhibit a close expression for the total number of interior points over all Motzkin polyominoes of length $n$.
@@ 2. Links between Motzkin words and other Motzkin classes
First, let us define recursively the bijection $\psi$ between ${\mathcal{M}}_{n}$ and the set of Motzkin paths of length $n-1$, i.e. lattice paths in the quarter plane, starting at the origin, ending on the $x$-axis, and made of $(n-1)$ steps $F=(1,0)$, $U=(1,1)$, and $D=(1,-1)$:
@@
where $(1+u)$ corresponds to the word obtained from $u$ by increasing by one all letters of $u$. For instance, we have
@@
On the other hand, in a previous paper [5], the authors study the avoidance of patterns in Catalan words of length $n$, i.e., words $w=w_{1}w_{2}\cdots w_{n}$ satisfying $w_{1}=\texttt{0}$ and $10\leq w_{i}\leq w_{i-1}+$ 1. They focus on Catalan words of length $n$ avoiding the pattern (,)($\geq$,$\geq$)( ≥ , ≥ ), that is those that do not contain occurrences $w_{i}\geq w_{i+1}\geq w_{i+2}$, or equivalently those avoiding the consecutive patterns $100$, $000$, $110$, and $210$. They prove algebraically that these words are counted by the Motzkin number $m_{n}$. Below, we give a bijection $\phi$ establishing a link between this class of Catalan words and ${\mathcal{M}}_{n}$.
Let $w$ be a Catalan word avoiding (,)($\geq$,$\geq$)( ≥ , ≥ ), then $\phi$ is recursively defined as follow.
@@
where $\bar{u}$ is obtained from $u$ by deleting the last symbol, and where $(1+u)$ corresponds to the word obtained from $u$ by increasing by one all letters of $u$. For instance, we have
@@
We leave as an exercise to check that $\phi$ is a bijection.
@@ 3. Area and semiperimeter statistics
By multiplying (4) by $v^{i-1}$ and summing over $i\geq 2$, and by considering (3), we obtain
@@
Simplifying the last expression, we obtain the desired result. ∎
@@ 4. The semiperimeter statistic
In order to compute $S(x,p):=A(x;p,1;1)$, we use the kernel method  (see [2, 24]). For this case the method consists in cancelling the coefficient of $A(x;p,1;v)$ by finding $v$ as an algebraic function $v_{0}$ of $x$. So, by taking $v=v_{0}$ such that $01-p^{2}xv_{0}+\frac{px}{1-v_{0}}=$ 0, namely
$v_{0}=\frac{1+p^{2}x-\sqrt{1-2p^{2}x-4p^{3}x^{2}+p^{4}x^{2}}}{2p^{2}x},$
we obtain the expression given in Theorem 4.1.
@@
Let us prove the statement by induction on $n$. The case $n=0$ is trivial since $\psi(\texttt{1})=\epsilon$. Let us assume the statement is true for $m\leq n$, and let us prove it for $n+1$. Let $w\in{\mathcal{M}}_{n+1}$ with semiperimeter $2(n+1)-k$. In the following of the proof, we denote by ${\texttt{up}}(P)$ the number of up-steps of the Motzkin path $M$.
If $w=\texttt{1}(1+u),u\neq\epsilon$, then we have ${\texttt{sper}}(u)=2n-k$, $\psi(w)=F\psi(u)$, and ${\texttt{up}}(\psi(w))={\texttt{up}}(\psi(u))$. Using the recurrence hypothesis on $u$, we have ${\texttt{up}}(\psi(u))=k$, and then ${\texttt{up}}(\psi(w))=k$ as expected.
If $w=\texttt{1}(1+u)v,u,v\neq\epsilon$, then we have ${\texttt{sper}}(u)=s_{1}$ and ${\texttt{sper}}(v)=s_{2}$ with $s_{1}+s_{2}+1=2(n+1)-k$, $\psi(w)=U\psi(u)D\psi(v)$, and ${\texttt{up}}(\psi(w))=1+{\texttt{up}}(\psi(u))+{\texttt{up}}(\psi(v))$. Using the recurrence hypothesis on $u$ and $v$, we have ${\texttt{up}}(\psi(u))=2|u|-s_{1}$, and ${\texttt{up}}(\psi(v))=2|v|-s_{2}$, where $|u|$ is the length of $u$. Hence, this implies that ${\texttt{up}}(\psi(w))=1+2|u|-s_{1}+2|v|-s_{2}=2+2|u|+2|v|-2(n+1)+k=k$, as expected.
Considering these two cases, the induction is completed ∎
@@
Let $s(n,i)$ denote the sum of the semiperimeter of all Motzkin polyominoes of length $n$ that last column has height $i$. The first few values are
@@
This sequence does not appear in the OIES. The expected value of the semiperimeter is $35n/$ 3. In Corollary 4.5 we give an explicit relation for the sequence $s(n)$ using the central trinomial coefficient $T_{n}$, that is the coefficient of $x^{n}$ in the expansion $(x^{2}+x+1)^{n}$. From the multinomial theorem we see that
@@
$\frac{\partial}{\partial p}S(x,p)\rvert_{q=1}$ $=\frac{1+x^{2}-(1+x)\sqrt{1-2x-3x^{2}}}{2x\sqrt{1-2x-3x^{2}}}$
@@ 5. The area statistic
$U(x,q)=\frac{\sum_{j\geq 1}(-1)^{j-1}q^{j}x^{j}\prod_{i=1}^{j-1}\frac{1-q^{i}+q^{2i}}{1-q^{i}}}{1-\sum_{j\geq 1}(-1)^{j-1}\frac{q^{j}x^{j}}{1-q^{j}v}\prod_{i=1}^{j-1}\frac{1-q^{i}+q^{2i}}{1-q^{i}}}.$
@@
First note that the Motzkin words of length $n$ are in bijection with Dyck paths of length $2n$ avoiding the subword $UDU$. Indeed, given a Dyck path, we associate a Motzkin word formed by the $y$-coordinate of each final point of the up steps. For example, in Figure 8 we show the Dyck path associated to the Motzkin word $\texttt{123212343}\in{\mathcal{M}}_{9}$. We denote by ${\mathcal{D}}$ the set of Dyck paths avoiding $UDU$.
@@
By differentiating twice this equation with respect to $v$ and taking $v=xM(x)+1$, we obtain that
@@ 5.1. Link with trinomial coefficients
The sequence $u(n)$ (A055217) corresponds to the sum of first $n$ coefficients of $(1+x+x^{2})^{n}$. Enigmatically, the monomials from the first part of the expansion of $(1+x+x^{2})^{n}$ can be literally written on the cells of all Motzkin polyominoes of size $n$ in a simple one-to-one manner. Term $x^{k}$ goes onto a cell of height $n-k$, see Figure 9. This subsection is devoted to explain this fact, which is easy to state, but not easy to prove. Theorem 5.3 follows from the results of this section if we replace $x$ by 1, consider the symmetry of coefficients of $(1+x+x^{2})^{n}$, and subtract the central trinomial coefficient.
@@
and $T(n,n-1)=T_{n-1}+T(n-1,n-2)+T(n-1,n-3)$, where $T_{n}$ is the central trinomial coefficient, that is $T_{n}=T(n,n)$. We also have $T(n,n-1)=n|{\mathcal{M}}_{n}|=nm_{n-1}$, see  A005717 in the OIES for more information. If $i>n-1$ or $i<0$ we define $T(n,i)=0$, moreover $T(1,1)=1$. The first few values of this array are
@@
Motzkin words of length $n$ are in bijection with Dyck paths of length $2n$ avoiding the subword $UDU$ (the $i$-th entry of the word $w$ corresponds to the $y$-coordinate of the endpoint of the $i$-th up-step of its corresponding Dyck path). From this bijection is clear that
@@
A grand Motzkin path of length $n$ is a Motzkin path without the condition that never passes below the $x$-axis. Let $g(n,i)$ be the number of grand Motzkin paths of length $n$ ending at height $i$. It is clear that $g(n,i)=g(n-1,i-1)+g(n-1,i)+g(n-1,i+1)$. We define the generating function
@@
Let $P$ be a grand Motzkin path ending at height $i$. It can be decomposed as $GUM_{1}UM_{2}U\cdots UM_{i}$, where $G$ is a grand Motzkin path ending at height zero and $M_{j}$ is a Motzkin path for all $1\leq j\leq i$. Therefore, we have the generating function $G_{i}(x)=x^{i}G_{0}(x)M^{i}(x)$, where $M(x)$ is the generating function of the Motzkin numbers. Notice that $G_{0}(x)=1+xG_{0}(x)+2x^{2}M(x)G_{0}(x)$, so $G_{0}(x)=1/(1-x-2x^{2}M(x))$. Therefore,
$G_{i}(x)=\frac{x^{i}M_{i}(x)}{1-x-2x^{2}M(x)}=B_{i}(x).$
@@
Second proof of Theorem 5.3. Notice that $\sum_{i=0}^{n-1}T(n,i)=a(n)$. But the trinomial coefficients are palindromic, that is $T(n,i)=T(n,2n-i)$, then
$\sum_{i=0}^{2n}T(n,i)=2a(n)-T(n,n)=2a(n)-T_{n}.$
Notice that we have to subtract the central trinomial coefficient. But it is clear that this sum is $3^{n}$, therefore $3^{n}=2a(n)-T_{n}$.
We end this section by asking the following (open) question: can we obtain a generalization of this result by using tetranomial coefficients (i.e. coefficients of the polynomials $(1+x+x^{2}+x^{3})^{n}$), and for which polyomino classes?

### v3

@@ Abstract.
A word $w=w_{1}\cdots w_{n}$ over the set of positive integers is a Motzkin word whenever $w_{1}=\texttt{1}$, $11\leq w_{k}\leq w_{k-1}+$ 1, and $w_{k-1}\neq w_{k}$ for $k=2,\dots,n$. It can be associated to a $n$-column Motzkin polyomino whose $i$-th column contains $w_{i}$ cells, and all columns are bottom-justified. We reveal bijective connections between Motzkin paths, restricted Catalan words, primitive Łukasiewicz paths, and Motzkin polyominoes. Using the aforementioned bijections together with classical one-to-one correspondence with Dyck paths avoiding $UDU$s, we provide generating functions with respect to the length, area, semiperimeter, value of the last symbol, and number of interior points of Motzkin polyominoes. We give asymptotics and closed-form expressions for the total area, total semiperimeter, sum of the last symbol values, and total number of interior points over all Motzkin polyominoes of a given length. We also present and prove an engaging trinomial relation concerning the number of cells lying at different levels and first terms of the expanded $(1+x+x^{2})^{n}$.
@@ 1. Introduction
These sequences have a similar behavior since, almost always, they appear together, and a large number of similar types of combinatorial classes are enumerated by these numbers. For Catalan numbers, we refer to the catalog of Catalan combinatorial classes compiled by Stanley [30]. Similarly, there are several combinatorial classes where the Motzkin family arises. For example, the rooted plane trees are a classical Catalan family, but if we add loops, yields a Motzkin family [14]. Additional examples of Motzkin families can be found in [4, 8] and the references therein.
@@
In this work, we focus on the class of Motzkin polyominoes, and we investigate enumerative problems about several statistics on these objects.
@@
Motivation and outline of the paper. The motivation of this work is to provide enumerative results on Motzkin words according to several parameters (area, semiperimeter, and number of interior points) defined on the associated polyominoes. Except for the first section where we exhibit bijections between Motzkin words and other classical combinatorial classes counted by the Motzkin numbers, all other results are obtained algebraically by considering functional equations for multivariate generating functions. In Section 2 we give three bijections between Motzkin words and three known combinatorial classes counted by the Motzkin numbers. In Section 3, we focus on the two statistics of the area and the semiperimeter. We provide the trivariate generating function, where the coefficient of $x^{n}p^{k}q^{\ell}$ is the number of Motzkin polyominoes $P(w)$ of length $n$, and satisfying ${\texttt{sper}}(w)=k$ and ${\texttt{area}}(w)=\ell$. From this, we deduce the generating function for the total sum of the last symbol over all Motzkin words of length $n$, and we give asymptotic approximation for the coefficient of $x^{n}$. In Section 4, we focus on the semiperimeter statistic. We give the generating function for nonempty Motzkin polyominoes with respect to the length and the semiperimeter. We exhibit a bijection between Motzkin polyominoes of length $n$ with semiperimeter $2n-k$, and Motzkin paths of length $n-1$ with exactly $k$ up steps. We deduce a closed-form expression for the total semiperimeter over all Motzkin polyominoes of length $n$. In Section 5, we make a similar study as in Section 4 by considering the area instead of the semiperimeter. Finally, Section 6 is dedicated to the statistic of the number of interior points of the Motzkin polyominoes, and we exhibit a closed-form expression for the total number of interior points over all Motzkin polyominoes of length $n$.
@@ 2. Links between Motzkin words and other Motzkin classes
First, let us recursively define the bijection $\psi$ between ${\mathcal{M}}_{n}$ and the set of Motzkin paths of length $n-1$, i.e. lattice paths in the first quadrant, starting at the origin, ending on the $x$-axis, and made of $(n-1)$ steps $F=(1,0)$, $U=(1,1)$, and $D=(1,-1)$:
@@
where $u$ and $v$ are Motzkin words and $(1+u)$ corresponds to the word obtained from $u$ by increasing by one all letters of $u$. For instance, we have
@@
On the other hand, in a previous paper [6], the authors study the avoidance of patterns in Catalan words of length $n$, i.e. words $w=w_{1}w_{2}\cdots w_{n}$ satisfying $w_{1}=\texttt{0}$ and $10\leq w_{i}\leq w_{i-1}+$ 1. They focus on Catalan words of length $n$ avoiding the pattern (,)($\geq$,$\geq$)( ≥ , ≥ ), that is those that do not contain occurrences $w_{i}\geq w_{i+1}\geq w_{i+2}$, or equivalently those avoiding the consecutive patterns $100$, $000$, $110$, and $210$. They prove algebraically that these words are counted by the Motzkin number $m_{n}$. Below, we give a bijection $\phi$ establishing a link between this class of Catalan words and ${\mathcal{M}}_{n}$.
Let $w$ be a Catalan word avoiding (,)($\geq$,$\geq$)( ≥ , ≥ ), then $\phi$ is recursively defined as follows.
@@
where $u$ and $v$ are some Catalan words, and $\bar{u}$ is obtained from $u$ by deleting the last symbol, and where $(1+u)$ corresponds to the word obtained from $u$ by increasing by one all letters of $u$. For instance, we have
@@
We leave it as an exercise to check that $\phi$ is a bijection.
@@ 3. Area and semiperimeter statistics
By multiplying equation (4) by $v^{i-1}$ and summing over $i\geq 2$, we obtain
@@
Simplifying the last expression using (3), we obtain the desired result. ∎
@@
The asymptotic approximation is easily obtained with a singularity analysis (see [15, 27]).
@@
Notice that we can retrieve Corollary 3.3 as follows. From a Motzkin word $w\in\mathcal{M}_{n-1,k}$, we can construct $k$ Motzkin words $w^{\prime}=wx\in\mathcal{M}_{n,x}$ by adding a letter $x\in[1,k-1]\backslash\{k\}$. Conversely, any Motzkin word of length $n$ can be uniquely constructed with this process. This means that $\sum_{k=1}^{n-1}km(n-1,k)=m_{n}$ which gives a simple combinatorial interpretation of Corollary 3.3.
@@ 4. The semiperimeter statistic
In order to compute $S(x,p):=A(x;p,1;1)$, we use the kernel method  (see [2, 28]). The method consists in cancelling the coefficient of $A(x;p,1;v)$ by taking the small root $v_{0}$ (the ones going to 00 for $x\sim 0$) of $1-p^{2}xv_{0}+\frac{px}{1-v_{0}}$, namely
$v_{0}=\frac{1+p^{2}x-\sqrt{1-2p^{2}x-4p^{3}x^{2}+p^{4}x^{2}}}{2p^{2}x}.$
When the factor on the left-hand side equals zero (and thus satisfies the equation), the right-hand side also equals zero and so the solution $v_{0}$ can be used on the right-hand side to obtain $A(x;p,1,1)$ (see Theorem 4.1).
@@
Let us prove the statement by induction on $n$. The case $n=0$ is trivial since $\psi(\texttt{1})=\epsilon$. Let us assume the statement is true for $m\leq n$, and let us prove it for $n+1$. Let $w\in{\mathcal{M}}_{n+1}$ with semiperimeter $2(n+1)-k$. We denote by ${\texttt{up}}(M)$ the number of up-steps of the Motzkin path $M$.
If $w=\texttt{1}(1+u),u\neq\epsilon$, then we have ${\texttt{sper}}(u)=2n-k$, $\psi(w)=F\psi(u)$, and ${\texttt{up}}(\psi(w))={\texttt{up}}(\psi(u))$. Using the induction hypothesis on $u$, we have ${\texttt{up}}(\psi(u))=k$, and then ${\texttt{up}}(\psi(w))=k$ as expected.
If $w=\texttt{1}(1+u)v,u,v\neq\epsilon$, then we have ${\texttt{sper}}(u)=s_{1}$ and ${\texttt{sper}}(v)=s_{2}$ with $s_{1}+s_{2}+1=2(n+1)-k$, $\psi(w)=U\psi(u)D\psi(v)$, and ${\texttt{up}}(\psi(w))=1+{\texttt{up}}(\psi(u))+{\texttt{up}}(\psi(v))$. Using the induction hypothesis on $u$ and $v$, we have ${\texttt{up}}(\psi(u))=2|u|-s_{1}$, and ${\texttt{up}}(\psi(v))=2|v|-s_{2}$, where $|u|$ is the length of $u$. Hence, this implies that ${\texttt{up}}(\psi(w))=1+2|u|-s_{1}+2|v|-s_{2}=2+2|u|+2|v|-2(n+1)+k=k$, as expected.
Considering these two cases, the induction is completed. ∎
@@
Let $s(n,i)$ denote the sum of the semiperimeters of all Motzkin polyominoes of length $n$ whose last column has height $i$. The first few values are
@@
Notice that from the decomposition given in Figure 5, we have for $n\geq 2$ and $2\leq i\leq n$,
@@
This sequence does not appear in the OEIS. The expected value of the semiperimeter is $35n/$ 3. In Corollary 4.5 we give an explicit relation for the sequence $s(n)$ using the central trinomial coefficient $T_{n}$, that is the coefficient of $x^{n}$ in the expansion $(x^{2}+x+1)^{n}$. From the multinomial theorem we see that
@@
$\frac{\partial}{\partial p}S(x,p)\rvert_{p=1}$ $=\frac{1+x^{2}-(1+x)\sqrt{1-2x-3x^{2}}}{2x\sqrt{1-2x-3x^{2}}}$
@@
We leave as an open question to find a combinatorial interpretation of the equation obtained in Corollary 4.5.
@@ 5. The area statistic
$U(x,q)=\frac{\sum_{j\geq 1}(-1)^{j-1}q^{j}x^{j}\prod_{i=1}^{j-1}\frac{1-q^{i}+q^{2i}}{1-q^{i}}}{1-\sum_{j\geq 1}(-1)^{j-1}\frac{q^{j}x^{j}}{1-q^{j}}\prod_{i=1}^{j-1}\frac{1-q^{i}+q^{2i}}{1-q^{i}}}.$
@@
First note that the Motzkin words of length $n$ are in bijection with Dyck paths of length $2n$ avoiding the consecutive subword $UDU$. Indeed, given a Dyck path avoiding $UDU$, we associate a Motzkin word formed by the $y$-coordinate of each final point of the up steps. For example, in Figure 8 we show the Dyck path associated to the Motzkin word $\texttt{123212343}\in{\mathcal{M}}_{9}$. We denote by ${\mathcal{D}}$ the set of Dyck paths avoiding $UDU$.
@@
By twice differentiating this equation with respect to $v$ and taking $v=xM(x)+1$, we obtain that
@@ 5.1. Link with trinomial coefficients
The sequence $u(n)$ (A055217) corresponds to the sum of the first $n$ coefficients of $(1+x+x^{2})^{n}$. Enigmatically, the monomials from the first part of the expansion of $(1+x+x^{2})^{n}$ can be literally written on the cells of all Motzkin polyominoes of size $n$ in a simple one-to-one manner. Term $x^{k}$ goes onto a cell of height $n-k$, see Figure 9. This subsection is devoted to explain this fact, which is easy to state, but not easy to prove. Theorem 5.3 follows from the results of this section if we replace $x$ by 1, consider the symmetry of coefficients of $(1+x+x^{2})^{n}$, and subtract the central trinomial coefficient.
@@
and $T(n,n-1)=T_{n-1}+T(n-1,n-2)+T(n-1,n-3)$, where $T_{n}$ is the central trinomial coefficient, that is $T_{n}=T(n,n)$. We also have $T(n,n-1)=n|{\mathcal{M}}_{n}|=nm_{n-1}$, see A005717 in the OEIS for more information. If $i>n-1$ or $i<0$ we define $T(n,i)=0$, moreover $T(1,1)=1$. The first few values of this array are
@@
Motzkin words of length $n$ are in bijection with Dyck paths of length $2n$ avoiding the subword $UDU$ (the $i$-th entry of the word $w$ corresponds to the $y$-coordinate of the endpoint of the $i$-th up-step of its corresponding Dyck path). From this bijection it is clear that
@@
A Motzkin walk of length $n$ is a Motzkin path prefix without the condition that never passes below the $x$-axis (see for instance [3]). A Motzkin walk ending on $x$-axis is also known as Grand Motzkin path. Let $g(n,i)$ be the number of Motzkin walks of length $n$ ending at height $i$. It is clear that $g(n,i)=g(n-1,i-1)+g(n-1,i)+g(n-1,i+1)$. We define the generating function
@@
Let $P$ be a Motzkin walk ending at height $i$. It can be decomposed as $GUM_{1}UM_{2}U\cdots UM_{i}$, where $G$ is a Motzkin walk ending at height zero and $M_{j}$ is a Motzkin path for all $1\leq j\leq i$. Therefore, we have the generating function $G_{i}(x)=x^{i}G_{0}(x)M^{i}(x)$, where $M(x)$ is the generating function of the Motzkin numbers. Notice that $G_{0}(x)=1+xG_{0}(x)+2x^{2}M(x)G_{0}(x)$, because a nonempty Motzkin walk ending at height zero is either $FG^{\prime}_{0}$, or $UMDG^{\prime}_{0}$, or $D\bar{M}UG^{\prime}_{0}$ where $G^{\prime}_{0}$ is a Motzkin walk ending at height zero, $M$ is a Motzkin path (ending on the $x$-axis) and $\bar{M}$ is a Motzkin walk obtained from a Motzkin path (ending on the $x$-axis) after exchanging $U$ and $D$. So, we have $G_{0}(x)=1/(1-x-2x^{2}M(x))$. Therefore,
$G_{i}(x)=\frac{x^{i}M^{i}(x)}{1-x-2x^{2}M(x)}=B_{i}(x).$
@@
Second proof of Theorem 5.3. Notice that $\sum_{i=0}^{n-1}T(n,i)=u(n)$. But the trinomial coefficients are palindromic, that is $T(n,i)=T(n,2n-i)$, then
$\sum_{i=0}^{2n}T(n,i)=2u(n)-T(n,n)=2u(n)-T_{n}.$
Notice that we have to subtract the central trinomial coefficient. But it is clear that this sum is $3^{n}$, therefore $3^{n}=2u(n)-T_{n}$.
We end this section by asking the following (open) questions: can we obtain an interpretation moving down a level to binomial coefficients? Can we obtain a generalization of this result by using tetranomial coefficients (i.e. coefficients of the polynomials $(1+x+x^{2}+x^{3})^{n}$), and for which polyomino classes?